The cost, in dollars, to produce x "I'd rather be a Sasquatch" T-Shirts is
C(x)=2x+26, x≥0 and the price-demand function, in dollars per shirt, is
p(x)= 30-2x, 0≤x≤15.
a)Find the profit function P(x).
b)Find the number of items which need to be sold in order to maximize profit.
c)Find the maximum profit.
d)Find the price to charge per item in order to maximize profit.
e)FIND AND INTERPRET THE BREAK EVEN POINTS.
I came up with the following solutions, which are correct per the answer section of my book:
a) P(x)= -2x²+28x-26
b) $7
c) Max profit: $72
d) price to charge for max. profit: $16
e) Cannot figure out. The given solutions per textbook are 1 and 13.
Please help me figure out:
FIND AND INTERPRET THE BREAK EVEN POINTS.
Thank you!!!
THANK YOU!

Answer by ewatrrr(24785)

e) Calculate and interpret the break-even point
the break-even point: R(x) = C(x)
(30x - 2x^2)=(2x+26)
2x^2 -28x + 26 = 0
x^2 - 14x + 13 = 0
(x-13)(x-1) = 0
x = 1, 13